# FA: Forced Prompt Learning of Vision-Language Models for Out-of-Distribution Detection

### v2

@@ 5.1 Experimental details
Datasets. We use a popular benchmark for conventional OOD detection [19, 30, 33, 32], where ImageNet-1k [7] serves as the ID dataset, and the OOD datasets are iNaturalist [17], SUN [47], Places [52], and Texture [6]. Moreover, inspired by existing studies [2, 1], we use datasets like OpenImage-O [45], NINCO [2], and ImageNet-O [15], which are cleaner and more realistic, to simulate the more challenging OOD detection. Besides, we also utilized other widely adopted datasets for few-shot settings as ID datasets [54, 49, 51], which include StandfordCars [21], UCF101 [40],Caltech101 [12], Flowers102 [35], EuroSAT [13], FGVCAircraft [29], OxfordPets [36], and Food101 [3], considering factors such as image resolution and the number of classes.
@@ 5.2 Main results
Challenging OOD Detection. Recent studies [2, 1] have shown that some of the used OOD datasets contain more or less images of objects that belong to ID classes in ImageNet-1k [7]. Hence, probing the true performance of OOD detectors for ImageNet-1k demands OOD datasets that are both challenging and genuinely OOD. As shown in Tab. 3, to fully substantiate the effectiveness of our FA, we leverage 3 additional cleaner and more challenging OOD datasets, including OpenImage-O [45], NINCO [2], and ImageNet-O [15], inspired by [2]. Notably, both our and consistently outperform current SOTA methods in terms of average FPR95 and AUROC under different few-shot settings (More details under the 1-shot and 4-shot setting are provided in the Appendix A).
@@ 7 Acknowledgments
Referring to Section 5.2, we also conduct the experiments under 4-shot setting on both conventional and challenging ImageNet-1k OOD benchmark. As shown in Tab. 8 and Tab. 9, our FA achieves superior average performance.

### v3

@@ 5.1 Experimental details
Datasets. We use a popular benchmark for conventional OOD detection [19, 30, 33, 32], where ImageNet-1k [7] serves as the ID dataset, and the OOD datasets are iNaturalist [17], SUN [47], Places [52], and Texture [6]. Moreover, inspired by existing studies [2, 1], we use datasets like OpenImage-O [45], NINCO [2], and ImageNet-O [15], which are cleaner and more realistic, to simulate the more challenging OOD detection. Besides, we also utilize other widely adopted datasets for few-shot settings as ID datasets [54, 49, 51], which include StandfordCars [21], UCF101 [40],Caltech101 [12], Flowers102 [35], EuroSAT [13], FGVCAircraft [29], OxfordPets [36], and Food101 [3], considering factors such as image resolution and the number of classes.
@@ 5.2 Main results
Challenging OOD Detection. Recent studies [2, 1] have shown that some of the used OOD datasets contain more or less images of objects that belong to ID classes in ImageNet-1k [7]. Hence, probing the true performance of OOD detectors for ImageNet-1k demands OOD datasets that are both challenging and genuinely OOD. As shown in Tab. 3, to fully substantiate the effectiveness of our FA, we leverage 3 additional cleaner and more challenging OOD datasets, including OpenImage-O [45], NINCO [2], and ImageNet-O [15], inspired by [2]. Notably, both our and consistently outperform current SOTA methods in terms of average FPR95 and AUROC under different few-shot settings (More details under the 1-shot and 4-shot settings are provided in the Appendix A).
@@ 7 Acknowledgments
Referring to Section 5.2, we also conduct experiments under other few-shot settings on both conventional and challenging ImageNet-1k OOD benchmarks. As shown in Tab. 8 and Tab. 9, our FA achieves superior average performance.